### v1

# Beyond Accuracy-Fairness: Stop evaluating bias mitigation methods solely on between-group metrics

Abstract
Artificial Intelligence (AI) finds widespread applications across various domains, sparking concerns about fairness in its deployment. While fairness in AI remains a central concern, the prevailing discourse often emphasizes outcome-based metrics without a nuanced consideration of the differential impacts within subgroups. Bias mitigation techniques do not only affect the ranking of pairs of instances across sensitive groups, but often also significantly affect the ranking of instances within these groups. Such changes are hard to explain and raise concerns regarding the validity of the intervention. Unfortunately, these effects largely remain under the radar in the accuracy-fairness evaluation framework that is usually applied. This paper challenges the prevailing metrics for assessing bias mitigation techniques, arguing that they do not take into account the changes within-groups and that the resulting prediction labels fall short of reflecting real-world scenarios. We propose a paradigm shift: initially, we should focus on generating the most precise ranking for each subgroup. Following this, individuals should be chosen from these rankings to meet both fairness standards and practical considerations.
Keywords Fairness, Transparency, Responsible AI, bias mitigation strategies
1 Introduction
In the rapidly evolving landscape of Artificial Intelligence (AI) and machine learning, the pursuit of fairness in algorithmic decision-making has emerged as a central concern. As the influence and scope of the decisions made by AI systems are increasing, there are growing concerns that the models making these decisions might unintentionally encode and even amplify human bias (Corbett-Davies and Goel, 2018). The term bias is used to describe the situation in which sensitive groups are substantially disadvantaged by an algorithm or model. One of the ways bias can seep into a model is when it is trained on biased data, following the famous garbage in, garbage out principle which emphasizes how flawed input data will result in flawed output (Geiger et al., 2020). Examples of bias in AI models are everywhere, with cases in almost every domain. In the context of hiring, a well-known use case is that of an automated Amazon recruitment system that had to be pulled because it was biased against female applicants (Dastin, 2022). Much earlier already, St George’s Hospital Medical School’s Commission for Racial Equality discovered that a computer program used for initial screenings of applicants “written after careful analysis of the way in which the staff were making these choices” unfairly rejected women and individuals with non-European sounding names (Johnson, 2021; Lowry and Macpherson, 1988). There is an abundance of examples akin to these ones.
Many bias detection and mitigation methods exist, but most of them focus on "between-group fairness" where the primary objective is to rectify disparities in model predictions between distinct demographic groups. This endeavor is undeniably critical, as it aims to rectify long-standing inequalities. However, it is equally essential to acknowledge and scrutinize the complexities that exist "within" these groups, giving rise to a concept that is commonly referred to as "within-group fairness". Speicher et al. (2018) already note that many approaches to group fairness tackle only between-group issues, worsening within-group fairness as a consequence. Krco et al. (2023) highlight that the blind optimization of commonly used fairness metrics does not show who is impacted within each group, while Mittelstadt et al. (2023) emphasize that many of the currently used bias mitigation methods can make every group worse off. These issues illustrate why solely looking at fairness by measuring disparities between groups is not adequate.
While benchmarking studies attempt to gauge the performance of various bias mitigation methods, they often fall short, comparing what can be described as ‘apples to oranges’. This issue arises because different bias mitigation methods can significantly vary the number of positive instances, and by comparing them as-is, the strict constraints faced by practitioners are not taken into account. In almost every application of machine learning, such constraints are a reality. For example, banks in credit scoring are limited by an annual loan quota. Universities have a cap on student admissions each year. In cancer screening, healthcare authorities can only screen a certain percentage of the population due to capacity constraints. Similarly, in predicting recidivism, the number of available prison places limits the potential positive rate.
The current state-of-art in benchmark studies of bias mitigation methods is to compare these methods on metrics based on the prediction label (such as accuracy or statistical parity difference) (Chen et al., 2023; Hort et al., 2022; Hufthammer et al., 2020; Reddy, 2022; Janssen and Sadowski, 2021; Menon and Williamson, 2018). However, the current metrics fall short in acknowledging the inherent differences in the situations they compare, rendering them insufficient for determining the superior performance of one method over another. More importantly, they fail to take into account the constraints in positive rate that are faced in the real world. For example, our results in Section 4 demonstrate that for the Adult Income dataset one bias mitigation method yields a positive rate of 0.5% while another results in a positive rate of 39.3%, whereas the actual positive rate of the dataset is 23.9%. If we compare the accuracy of these two mitigation methods, we are comparing two entirely distinct points of the data distribution, both of which may be inapplicable in real-world settings. Moreover, the majority of bias mitigation methods do not result in the satisfaction of the fairness metrics, necessitating additional post-hoc interventions after their application.
We argue that the focus should shift to fairness within groups, and the crucial question to be answered should be: how can we select the best individuals within each group? Machine learning models not only return prediction labels, they also return prediction scores, which creates an inherent ranking of instances. Given that the prediction labels from many mitigation methods are often impractical due to their unrealistic positive rates and non-compliance with the selected fairness metric, it is more logical to concentrate on the prediction scores produced by each mitigation method. The objective should be to have the machine learning model rank each subgroup as accurately as possible, a metric conveniently measured by the Area Under the Curve (AUC). Once the most accurate ranking is established, post-hoc fairness methods can be employed to select the highest-ranked individuals within each group, aligning with both industry constraints and fairness metrics. [^1] This framework operates under the assumption of having access to the sensitive attribute(s).
Figure 1: The ranking within each group is shifted by most preprocessing and inprocessing bias mitigation methods. The current state-of-the-art is to measure the accuracy and fairness metrics at one specific classification threshold.
When striving to establish the most optimal ranking, it is crucial to differentiate between pre- and inprocessing bias mitigation methods on the one hand, and postprocessing methods on the other. Only pre- and inprocessing methods have the capacity to alter this ranking by either fine-tuning the training data or incorporating fairness constraints during model training, as this will change the prediction scores of the machine learning model. In contrast, postprocessing methods merely flip prediction labels based on the scores of the initial machine learning model, without altering the inherent ranking of instances. Figure 1 visualizes how the inherent ranking is shifted by pre- and inprocessing mitigation methods, emphasizing the prevailing practice of assessing performance and fairness metrics at a single classification threshold. Furthermore, it underscores the tendency to overlook the impact of fairness interventions on the rankings within each subgroup.
The contributions of this paper are threefold: Firstly, we provide insights into the operational dynamics of bias mitigation methods and illustrate how some methods will significantly impact the ranking within each group, while others will not. Secondly, we criticize the current approach to comparing bias mitigation methods, as it compares situations that are significantly distinct and not applicable in real-world settings. Therefore, we propose a paradigm shift, advocating for a transition from focusing on measuring the disparity in prediction labels between-groups to establishing the best prediction scores within-groups. This leads us to our last contribution: a novel framework. This framework consists of a two-step approach: first, optimizing for the most accurate ranking within each subgroup, and then, selecting the highest-ranked individuals from each subgroup, aligning with fairness metrics and industry constraints.
2 Background
Before we go further, it is important to define some of the key terminology that is often used in the fairness literature. A sensitive attribute refers to a characteristic or feature(s) of an individual that are considered sensitive, often with respect to potential discrimination. This can include attributes such as race, gender, age, religion, sexual orientation, or any other factor that could be the basis of unfair treatment. Consequently, a protected group typically refers to the demographic group that is at risk of being unfairly treated or discriminated against based on their sensitive attribute, while the privileged group is the demographic category that is typically not subject to unfair treatment based on that sensitive attribute. In this paper, we operate under the assumption of a single binary sensitive attribute, implying the existence of a protected group and a privileged group. Fairness metrics are quantitative measures used to assess the fairness of AI models, while fairness (or bias) mitigation strategies are techniques and approaches used to modify AI models to reduce bias or discrimination.
In the computer science community, a plethora of fairness metrics have been proposed (Corbett-Davies and Goel, 2018). One of the most popular approaches is the group fairness metric, which quantifies the fairness of a machine learning model across different demographic or sensitive groups, aiming to identify disparities in the outcome between these groups. One of the simplest and most commonly used definitions is demographic parity (or statistical parity), which states that the positive classification rate must be the same regardless of the value of the protected attribute. In our example of hiring, this means that a model must invite equal percentages of white and black applicants for an interview (if race is the sensitive attribute) or of male and female applicants (if gender is the sensitive attribute). Other commonly used metrics include equalized opportunity, which states that there should be an equal proportion of true positives in both groups, and equalized odds, which examines whether both the proportion of true positives and true negatives is equal across groups. Besides these, many other fairness metrics exist, and the issue is that most of them cannot be optimized at the same time (Kleinberg et al., 2016). Deciding upon a group fairness metric to optimize thus means already imposing a certain world view.
A common starting point for designing a fair algorithm is simply to exclude sensitive attributes from the model, which is known as fairness through unawareness. The limitations of fairness through unawareness have been commonly addressed, with the most fundamental limitation being ‘the proxy problem’. The proxy problem states that the omission of sensitive attributes can lead the machine learning model to rely on proxy variables that indirectly encode the information contained in the sensitive attribute and hence still introduce bias into the model’s decision-making process. A classic example of the proxy problem is the use of zip codes in the United States as a proxy for racial information, as these two attributes tend to be heavily correlated. This has prompted many to argue that proxies should be excluded from the dataset as well, however, this is very difficult to operationalize (Corbett-Davies and Goel, 2018). This is because every attribute used in the machine learning model can be at least partially correlated with the sensitive attribute; and often, even strongly correlated attributes may be considered legitimate factors on which to base decisions (for example, education in the case of hiring) (Corbett-Davies and Goel, 2018).
This illustrates that creating a fair machine learning model is a tedious process. In response, many bias mitigation methods that claim to improve fairness, have been introduced. We can divide most of them into 3 categories: preprocessing, inprocessing and postprocessing. The idea behind preprocessing methods is that they will change the representation of the data before the machine learning model is learned. Inprocessing methods improve fairness during the training process, while postprocessing methods attempt to modify the posteriors in a way that satisfies fairness constraints. These bias mitigation methods focus on satisfying the aforementioned group fairness metrics that measure disparities between groups (Chen et al., 2023).
3 Materials and methods
3.1 Materials
We will use several real world datasets that are common in the domain of fair machine learning (Le Quy et al., 2022). The Adult Income dataset contains information extracted from the 1994 census data with as target variable whether the income of a person exceeds \$50,000 a year or not. The Compas dataset includes demographic information and criminal histories of defendants from Broward County, and is used to predict whether a defendant will re-offend within two years. The Dutch Census dataset represents aggregated groups of people in the Netherlands for the year 2001, and can be used to predict whether a person’s occupation can be categorized as a high-level (prestigious) or a low-level profession (Van der Laan, 2000). The Law admission dataset contains a Law School Admission Council (LSAC) survey conducted across 163 law schools in the United States in 1991 (Wightman, 1998) and predicts whether the student will pass the bar exam or not. The Student Performance dataset describes the achievements of students in two Portuguese schools (Cortez and Silva, 2008). The classification task is to predict whether they score above average in mathematics.
Table 1: Used datasets
Name # instances # attributes Protected attribute Protected group Target attribute Positive rate Adult 48,842 10 Gender Female High income 23.93% Compas 5,278 7 Race African-American Low risk 52.16% Dutch Census 60,420 11 Gender Female High occupation 52.39% Law admission 20,798 11 Race Non-White Pass the bar 88.97% Student Performance 649 29 Gender Male High score in mathematics 53.62%
3.2 Methods
3.2.1 Machine learning models
We train a fully connected neural network on each dataset, utilizing binary cross-entropy as the loss metric. [^2] We use a state-of-the-art machine learning model, with one hidden layer with 200 nodes. We use 50 epochs, a batch size of 128 and the Adam optimizer. We do not perform hyperparameter tuning. The choice of neural networks is driven by the necessity to accommodate one of our bias mitigation methods, Adversarial Debiasing, which exclusively operates with this type of model. Although the remaining methods are model-agnostic, we opt for consistency in our approach, employing a neural network across all bias mitigation methods to ensure comparability in our final results.
3.2.2 Bias mitigation methods
Numerous debiasing strategies currently exist, but we focus on the methods available in the AIF360 package (Bellamy et al., 2018). For all algorithms, we use the default settings.
As preprocessing methods, we use Learning Fair Representations (LFR) and Disparate Impact Remover (DIR). The idea behind Learning Fair Representations (Zemel et al., 2013) is that a new representation Z is learned that removes the information correlation with the sensitive attribute, but preserves the other information about X as much as possible. Disparate Impact Remover (Feldman et al., 2015) modifies the training data to reduce the influence of sensitive attributes, but preserves rank-ordering within groups.
We use the inprocessing methods of Adversarial Debiasing (ADV) and the Meta Fair Classifier (MFC). Adversarial Debiasing (Zhang et al., 2018) combines a classifier that predicts the class label with an adversary that predicts the sensitive attribute. The goal is to maximize the classifier’s performance while minimizing that of the adversary. The Meta Fair Classifier (Celis et al., 2019) takes the fairness metric as part of the input and returns a classifier optimized with respect to that fairness metric.
As postprocessing methods, we use Equalized Odds Postprocessing (EQP), Reject Option Classification (ROC) and Threshold Optimation (TO). Equalized Odds Postprocessing (Hardt et al., 2016) will solve a linear program to find the probabilities with which to change output labels in order to optimize equalized odds, while Reject Option Classification (Kamiran et al., 2012) will flip the predictions the model is not confident of. [^3] Reject option classification identifies instances where the model is uncertain about its prediction and essentially ‘rejects’ making a definite decision. In this implementation, aiming to enhance fairness, the labels of these instances are flipped to satisfy a fairness criteria. Threshold Optimization (Kamiran et al., 2012) is maybe the most straightforward method of mitigating bias as it will optimizes the thresholds for both groups in isolation. [^4] For this bias mitigation method, we use the implementation available through the Fairlearn toolbox (Bird et al., 2020) as this method is not available in the AIF360 toolbox. ROC and TO are implemented to enforce Demographic Parity, while EOP enforces Equalized Odds by default.
3.3 Metrics
3.3.1 Performance metrics
Most benchmark studies compare the different mitigation methods on accuracy, which measures how often the prediction label assigned by the machine learning model coincides with the true label (Chen et al., 2023; Hort et al., 2021; Krco et al., 2023).
It is commonly acknowledged that accuracy is not always the best metric to measure the performance of a machine learning model, as it is for example not suitable to deal with imbalanced class distributions (as in this case a model can obtain a high accuracy by just predicting all samples as the majority class) (Chen et al., 2023; Mittelstadt et al., 2023). This has led some studies to include other metrics such as the F1-score, Precision, or Recall (Chen et al., 2023). However, another notable drawback of these performance metrics is that they measure the performance at a specific classification threshold, as they use the prediction labels of the machine learning model and not the prediction scores. We can evaluate the performance of the prediction scores by using the Area Under the ROC Curve (AUC).
AUC allows for an objective comparison across classifiers, as it is unaffected by the choice of threshold or the frequency of classes (Hand, 2009). It measures how well the prediction scores (and thus the ranking) of a machine learning model distinguishes between positive and negative cases. The formula for the AUC score:
$P(S(x_{i})>S(x_{z})|y_{i}=1,y_{z}=0)$
This formula means that the AUC score is equivalent to the probability that a classifier will rank a randomly chosen positive instance higher than a randomly chosen negative instance. Provost et al. (1998) previously advocated for the adoption of AUC as a standard for comparing classifiers in the broader field of machine learning. Despite this, its integration into Fair Machine Learning has been limited. Furthermore, the specific context of Fair Machine Learning presents additional justifications for focusing on the prediction scores generated by machine learning models, rather than solely on their prediction labels, due to unrealistic positive rates and the unsatisfaction of required fairness metrics.
In this study, we operate under the assumption that label bias is absent, meaning that the actual labels accurately represent the intended prediction target (Wick et al., 2019; Favier et al., 2023; Lenders and Calders, 2023). Note that if label bias was present, it would compromise the validity of both the AUC and accuracy metrics, as these measures rely on these labels for their calculations.
Lastly, we will measure the positive classification rate of the whole population. As mentioned earlier, this is important to be able to compare the different bias mitigation methods. The formula for the Positive Rate (PR):
$P(\hat{Y}=1)$
3.3.2 Fairness metrics
Many possible fairness metrics exist, but we will report two metrics that are commonly used in the fairness domain to measure disparities between groups: Demographic parity (or statistical parity) states that the positive classification rate must be approximately the same in the protected group as in the privileged group. Statistical Parity Difference measures the difference in this positive rate between both groups. The formula for the Statistical Parity Difference (SPD):
$P(\hat{Y}=1|S=s)-P(\hat{Y}=1|S=ns)$
Equalized opportunity requires the true positive rate to be approximately the same across groups. The formula for Equal Opportunity Difference (EOD):
$P(\hat{Y}=1|S=s,Y=1)-P(\hat{Y}=1|S=ns,Y=1)$
Larger values of these metrics correspond to a higher level of bias towards one of the sensitive groups (Hort et al., 2021).
4 Results
4.1 Whom do the bias mitigation strategies affect?
The lack of transparency in the field of fair machine learning has been acknowledged in literature (Rudin et al., 2020; Wachter et al., 2021; Goethals et al., 2023), while some have already specifically criticized the opacity in the exact effects of the different bias mitigation methods (Krco et al., 2023; Favier et al., 2023; Holstein et al., 2019). We aim to shed light on the operational dynamics of the different bias mitigation methods, by comparing the score distributions after deploying each bias mitigation method with the score distributions of the initial machine learning model. We illustrate the prediction scores from the initial ML model on the x-axis and the scores post-application of various bias mitigation methods on the y-axis. Additionally, we categorize the instances based on their affiliation with either the protected group (represented in dark blue) or the privileged group (represented in light blue). The chart is divided into four quadrants, each depicting the classification of instances as either positive or negative by the initial ML model and by each bias mitigation method. The presented results in Figure 2 are calculated using the Compas dataset. The figures for the other datasets can be found in the Appendix (Figures 5, 6, 7 and 8), with the results being consistent to the results of the Compas dataset.
(a) ML model
(b) Preprocessing
(c) Preprocessing
(d) Inprocessing
(e) Inprocessing
(f) Postprocessing
(g) Postprocessing
(h) Postprocessing
Figure 2: Score distributions for the Compas dataset. The x-axis represents the prediction scores of the initial ML model, while the y-axis represents the prediction score after applying each bias mitigation method.
In Figure 1(a), we observe the prediction scores of the initial ML model. By default, machine learning models use a classification threshold of $0.5$, categorizing instances above this threshold as positive and those below it as negative. This behavior remains consistent for both the protected and privileged groups.
In Figures 1(b)-1(e), we can assess the relation between the prediction scores of the initial ML model and the prediction scores after applying preprocessing (LFR and DIR) or inprocessing (ADV and MFC) bias mitigation methods. Figure 1(b) reveals that LFR significantly alters the prediction scores, in a seemingly random way. Instances that receive a positive label post-LFR are notably different from those in the initial score ranking. This suggests a substantial transformation in the feature space due to the preprocessing method. We will assess how this affects the performance and fairness metrics in Section 4.2. As shown in Figures 1(c), 1(d), and 1(e), the methods DIR, ADV, and MFC also introduce significant alterations to the prediction scores, and thus the inherent ranking of instances. However, the altered prediction scores align more closely with the initial prediction scores compared to those obtained using LFR.
Figure 2 also provides us with insights into the fundamental distinction among various bias mitigation methods—specifically, whether they impact the prediction scores or only the prediction labels. Unlike preprocessing and inprocessing methods, postprocessing methods do not alter the inherent ranking; instead, they adjust labels based on the prediction scores of the initial machine learning model. Each method implements a unique strategy for label flipping, all with the objective of meeting specified fairness metrics. For instance, ROC (Figure 1(f)) targets the most uncertain instances for label flipping, TO (Figure 1(h)) establishes distinct thresholds for each group, while EOP (Figure 1(g)) determines the ideal quantity of labels to flip within each group and executes these flips randomly. In the scenarios of ROC and TO, group-specific new thresholds are determined (as illustrated in Figures  1(f) and 1(h)). Regarding EOP, label flipping occurs randomly among negatively classified individuals in the protected group and positively classified individuals in the privileged group (as shown in green and red in Figure 1(g)). In this case, the initial ranking impacts the initial labels, but will have no influence on the decision regarding which labels are flipped.
Our examination provides insights into the specific instances affected by each bias mitigation method, in response to the work of Krco et al. (2023). Take, for instance, the case of ROC: the instances that are flipped from a negative classification to a positive classification, are those instances that were initially classified as negative but that posses the highest prediction score according to the initial machine learning model.
Table 2: Table with the similarity between the initial ranking produced by the machine learning model and the ranking produced by the model after using a bias mitigation method, measured by the Kendall-Tau statistic. The values between brackets present the similarity in ranking for the protected group and privileged group respectively. All values are between zero and one, and higher values point to more similarity in ranking.
Dataset LFR DIR ADV MFC ROC-EOP-TO
Adult
0.318
(0.002, 0.034)
0.770
(0.726, 0.761)
0.729
(0.780,0.812)
0.638
(0.640, 0.663)
1
Compas
0.408
(0.153,0.534)
0.917
(0.883, 0.960)
0.805
(0.926, 0.946)
0.909
(0.900, 0.935)
1
Dutch
0.361
(-0.007, 0-0.004)
0.992
(0.991, 0.991)
0.808
(0.812, 0.913)
0.819
(0.850, 0.786)
1
Law
0.570
(0.622, 0.523)
0.870
(0.882, 0.858)
0.908
(0.896, 0.907)
-0.758
(-0.872, -0.737)
1
Student
0.215
(0.248, 0.216)
0.890
(0.891, 0.891)
0.515
(0.731, 0.678)
0.775
(0.843, 0.794)
1
Instead of visualizing how the prediction scores of the initial ML relate to the prediction scores after each bias mitigation method, we can also calculate the overlap between the two rankings by utilizing the Kendall-Tau statistic (Kendall, 1948). The outcomes are presented in Table 2. Table 2 illustrates that LFR consistently presents the least similarity among rankings. In contrast, the use of DIR leads to a significantly higher degree of overlap in rankings across all datasets. The similarity in rankings when applying ADV and MFC falls into a more moderate category. These observations highlight the varying degrees of impact each method exerts on the prediction scores and thus the ranking of individuals. With postprocessing methods, the ranking remains identical to the ranking of the initial model, resulting in a ranking similarity of 1 (refer to Table 2), and the AUC remains consistent with the initial model (see Table 3). Our analysis reveals that across all datasets, the prediction scores and consequently the intrinsic rankings are substantially modified by every preprocessing and inprocessing bias mitigation method we evaluated. While such modifications are not necessarily negative, it is crucial to assess whether they enhance or worsen the ranking, and to what extent. The specifics of these findings will be presented in Section 4.2. We propose that for the dual objectives of fairness and accuracy optimization, preprocessing and inprocessing mitigation methods should be adopted only if they improve subgroup rankings. Should they fail to do so, utilizing the ranking produced by the original machine learning model and then applying fairness interventions post-hoc may be a more effective strategy.
4.2 Can we compare the bias mitigation methods?
Table 3: Results of the bias mitigation strategies on the five datasets. We report: the AUC score overall, and split over the protected group and the privileged group, the accuracy (ACC), statical parity difference (SPD), equal opportunity difference (EOD) and the positive rate (PR). Best values are highlighted in bold.
Dataset Metric ML model LFR DIR ADV MFC ROC EOP TO Adult AUC 0.843 0.623 0.85 0.847 0.826 0.843 0.843 0.843 $AUC^{pro},AUC^{pri}$ 0.811, 0.826 0.501, 0.508 0.824, 0.835 0.834, 0.843 0.789, 0.816 0.811, 0.826 0.811, 0.826 0.811, 0.826 ACC 0.806 0.766 0.825 0.82 0.81 0.728 0.77 0.793 SPD -0.26 -0.005 -0.164 -0.09 -0.139 -0.055 -0.046 -0.005 EOD -0.139 0.0 -0.181 -0.274 -0.278 -0.332 -0.283 -0.368 PR 0.247 0.005 0.174 0.161 0.178 0.393 0.151 0.167 Compas AUC 0.834 0.693 0.833 0.808 0.832 0.834 0.834 0.834 $AUC^{pro},AUC^{pri}$ 0.814, 0.821 0.588, 0.703 0.81, 0.823 0.809, 0.815 0.812, 0.821 0.814, 0.821 0.814, 0.821 0.814, 0.821 ACC 0.758 0.645 0.753 0.737 0.736 0.732 0.683 0.727 SPD -0.376 -0.866 -0.332 -0.064 -0.211 -0.031 -0.07 0.014 EOD -0.13 -0.055 -0.131 -0.231 -0.175 -0.233 -0.25 -0.262 PR 0.518 0.49 0.617 0.58 0.696 0.528 0.545 0.569 Dutch AUC 0.887 0.657 0.887 0.874 0.883 0.887 0.887 0.887 $AUC^{pro},AUC^{pri}$ 0.884, 0.848 0.499, 0.498 0.884, 0.849 0.881, 0.847 0.89, 0.852 0.884, 0.848 0.884, 0.848 0.884, 0.848 ACC 0.812 0.476 0.786 0.768 0.695 0.776 0.754 0.763 SPD -0.318 0.0 -0.432 -0.171 -0.394 -0.066 -0.159 -0.02 EOD -0.026 -0.315 -0.047 -0.073 -0.024 -0.217 -0.243 -0.254 PR 0.416 1.0 0.586 0.303 0.203 0.395 0.45 0.396 Law AUC 0.882 0.83 0.883 0.879 0.122 0.882 0.882 0.882 $AUC^{pro},AUC^{pri}$ 0.848, 0.864 0.803, 0.792 0.853, 0.864 0.843, 0.862 0.146, 0.142 0.848, 0.864 0.848, 0.864 0.848, 0.864 ACC 0.903 0.897 0.903 0.901 0.22 0.772 0.879 0.892 SPD -0.197 -0.207 -0.184 -0.141 0.494 -0.044 -0.021 -0.002 EOD -0.111 -0.128 -0.122 -0.143 -0.151 -0.104 -0.197 -0.21 PR 0.954 0.967 0.961 0.965 0.289 0.711 0.961 0.979 Student AUC 0.803 0.693 0.817 0.772 0.762 0.803 0.803 0.803 $AUC^{pro},AUC^{pri}$ 0.819, 0.788 0.689, 0.642 0.819, 0.814 0.757, 0.779 0.781, 0.794 0.819, 0.788 0.819, 0.788 0.819, 0.788 ACC 0.759 0.626 0.759 0.703 0.738 0.697 0.728 0.728 SPD -0.104 -0.72 -0.065 -0.573 0.031 0.076 -0.016 0.071 EOD -0.166 -0.063 -0.202 0.027 -0.283 -0.191 -0.205 -0.246 PR 0.585 0.697 0.574 0.662 0.677 0.492 0.595 0.605
We present the results of performance and fairness metrics for bias mitigation methods across five datasets in Table 3.
Aligning with existing literature, we observe that fairness metrics such as SPD and EOD often yield conflicting results (Kleinberg et al., 2016). It is rarely the same method that returns the best results for both metrics. Furthermore, no single method consistently outperforms all others for one of the metrics. However, DIR stands out for its excellent AUC performance, aligning with its design to preserve rank-ordering within groups (Feldman et al., 2015).
When comparing bias mitigation methods, conventional benchmarking studies often focus on accuracy and fairness (Chen et al., 2023; Zemel et al., 2013). However, relying solely on accuracy, which uses prediction labels at a specific threshold, may lead to inappropriate comparisons. There are two reasons why using the prediction labels is not a suitable approach: First of all, the fairness metrics are not yet satisfied, and secondly, different mitigation methods can result in varying positive rates. Both reasons would lead to an additional altering of the prediction labels post-hoc, so comparing them at this stage does not seem sensible. A more comprehensive approach involves assessing the performance of the prediction scores.
Addressing the first concern, we see that in the large majority of cases, bias mitigation methods fail to satisfy a fairness metric, which is confirmed in literature (Chen et al., 2023). Particularly, only postprocessing strategies tend to show a high success rate in achieving the optimized fairness metric, in contrast to the preprocessing and inprocessing methods. As highlighted in other benchmark studies (Chen et al., 2023), we also note that employing bias mitigation methods can even lead to situations that are more unfair in terms of disparities between groups.
Regarding the second concern, Table 3 reveals significant positive rate disparities among different strategies. This inconsistency poses challenges in real-world applications, where a fixed or reasonably bounded positive rate is typically expected. For example, when attempting to satisfy one of the fairness metrics, practitioners might consider LFR in the Adult Income dataset, despite a slight accuracy loss. However, this choice results in an unexpectedly low positive rate (0.5%), deviating significantly from other methods. Similarly, for the Dutch dataset, MFC leads to the best value for EOD, but it only has a positive rate of 20.3%, while the initial model has a positive rate of 41.6%. DIR leads to the best AUC, but has a positive rate of 58.6%. We argue against treating methods with significantly different positive rates as comparable situations. In practice, most real-world applications will have a relatively fixed positive rate, and bias mitigation methods must be adapted accordingly.
Both these concerns can effectively be addressed by adjusting the classification threshold(s) of the predictive model. Consequently, it is more logical to evaluate these mitigation methods based on their prediction scores, rather than on their prediction labels, which are still subject to change. We advocate to compare these mitigation methods based on the AUC score, to assess how well the indiduals are ranked within each group, and to generate the prediction labels post-hoc, based on the chosen fairness metric and practical constraints.
It is crucial not only to examine the overall AUC score across the entire population, but also to scrutinize the AUC scores disaggregated by different groups. Certain bias mitigation methods may yield the optimal ranking for one group but not for another. This prompts the need for decision-making: should preference be given to the overall best ranking or to narrowing the gap in rankings between the protected and privileged groups? Alternatively, deploying two separate models could be considered to ensure the best ranking for each subgroup. Given these results, one should evaluate whether a slight improvement in AUC justifies the adoption of distinct models. Additionally, note that for both the Adult and the Dutch datasets, the LFR method reduces the within-group ranking to an essentially random ordering. In any scenario, deploying a bias mitigation method that decrease the AUC score for every subgroup is counterproductive when we want to optimize for both performance and fairness. [^5] Note that we operate under the assumption that there is no label bias. Unfortunately, this is a common outcome in practice, as many bias mitigation methods may compromise AUC scores across subgroups in an attempt to adhere to a fairness metric (Mittelstadt et al., 2023).
Note that the goal of this comparison is not to declare the superior performance of one of the mitigation methods. For this, a more comprehensive benchmark study is needed with a larger number of datasets, machine learning models, and extensive tuning of each bias mitigation method. This was already the goal of multiple other benchmark studies (Chen et al., 2023; Hort et al., 2021; Reddy, 2022). Our primary goal was to emphasize the inadequacy of the current way of benchmarking these methods with each other, based on prediction labels that should still be subject to change.
4.3 How should bias mitigation methods be implemented?
Figure 3: Proposed framework to mitigate bias in machine learning models. Note that all actions mentioned under Step 1 are optional, and can be used in combination with each other.
We propose a novel framework for implementing bias mitigation strategies, outlined in Figure 3. In the initial step, the focus should be on optimizing the ranking accuracy for each subgroup independently. Evaluating the AUC score for each subgroup is crucial, as a lower AUC score for one subgroup can arise from various factors. For example, there may be insufficient data available for this subgroup, limiting the model’s ability to accurately capture its characteristics. Additionally, the data quality for the protected group could be notably lower, introducing noise or inaccuracies into the learning process. Moreover, the model might learn patterns based on historical inequalities that are currently undergoing changes. The evolving nature of societal dynamics can lead to shifts in behavior and characteristics within subgroups, impacting the model’s predictive accuracy. Another potential reason is that the model learned patterns primarily from the privileged group, often the majority group, which do not accurately resemble the behavior of the protected group. Whatever the cause, the first step should be to enhance the ranking accuracy within each group. This can be done by selecting more data from that subgroup, improving data quality, employing feature engineering, deploying separate models for each group, removing sensitive attributes, or utilizing preprocessing or inprocessing mitigation techniques. Note that all these proposed interventions are optional, as they can also lead to a lower AUC score. This list is by no means exhaustive, and a combination of multiple strategies may be the most effective.
Subsequently, the second step involves selecting the top-ranked individuals from each group, aligning with the desired fairness metrics and the applicable industry constraints. For instance, in the context of college admissions, adhering to demographic parity and having the capacity to admit approximately 10% of applicants could entail selecting the 10% highest-ranked instances from each subgroup. This two-step framework aims to not only enhance the accuracy of rankings within subgroups but also align the selection process with predefined fairness criteria and practical considerations.
Figure 4: Accuracy when selecting the top ranked individuals using the prediction scores produced by MFC and the initial ML model. This analysis is done for for different positive rates.
Let’s walk through the application of our framework using the Dutch dataset. Table 3 reveals that the best ranking results for each subgroup are achieved with the MFC method. However, this improvement comes at the cost of a significant decrease in accuracy, dropping from 0.812 to 0.695. It’s important to note that comparing these situations directly may not be entirely fair, given that the positive rate of the initial machine learning model is more than double than the positive rate after applying MFC. Additionally, in both cases, the fairness metrics are not satisfied.
Now, let’s apply our framework and consider enforcing demographic parity. We select the top-ranked individuals from each group based on the prediction scores after applying MFC, and for the sake of comparison also after applying the initial machine learning model. We illustrate the resulting accuracy for various positive rates in Figure 4. Notably, for all tested positive rates, the accuracy achieved using the scores produced by MFC is as high as or slightly higher than the scores from the initial machine learning model.
Two key points should be highlighted in interpreting these results. First, we refrain from claiming that the ranking produced by MFC is optimal. Achieving optimality would require testing more machine learning models, finetuning all ML models and mitigation methods, or testing some of the other suggestions made in Figure 3. Our aim is to underscore that relying solely on accuracy metrics can lead to misinterpretations, as our subsequent steps should use the prediction scores rather than the prediction labels.
Additionally, it’s crucial to acknowledge that in this specific case, we selected demographic parity as the fairness metric, which is easily enforced by selecting the top-ranked individuals from each subgroup. Other fairness metrics may require using postprocessing mitigation methods or involve solving an optimization problem that considers both the fairness metric and the desired positive rate.
5 Discussion
In this paper, our initial aim is to provide clarity on the specific impact of bias mitigation methods on various instances. A crucial aspect involves comparing the ranking of instances before and after the application of bias mitigation methods. We observe that preprocessing and inprocessing methods introduce changes to the ranking, whereas postprocessing methods maintain the original ranking and simply adjust labels based on it.
Furthermore, we contend that the current approach to comparing bias mitigation methods is insufficient due to the sole focus on between-group metrics and the comparison of disparate scenarios. To address this limitation, we propose a novel evaluation framework. The framework comprises two key steps: Firstly, establish the most accurate ranking for each subgroup. Secondly, select the highest-ranked individuals from each subgroup, guided by fairness metrics and industry-specific constraints. This refined framework aims to provide a more nuanced and comprehensive assessment of bias mitigation methods, aligning with the intricacies of real-world applications. This aligns with a paradigm shift from between-group fairness to within-group fairness, where the goal is to do as well as possible for each subgroup. By incorporating the existing capacity in positive instances as a constraint within our framework, we effectively prevent scenarios where the positive rate is excessively high and unrealistic, as well as situations where capacity remains underutilized. Regrettably, this underutilization is a frequent outcome in the application of bias mitigation strategies, where the positive rate of the privileged group is often reduced to match that of the underprivileged group, resulting in no significant improvement for either group (Mittelstadt et al., 2023).
Although we posit that for optimal fairness and performance, pre- and inprocessing methods should only be employed if they enhance subgroup rankings, there are scenarios where employing these methods might be beneficial, even if they result in less favorable subgroup rankings. Our framework can lead to situations where two otherwise identical entities are treated differently based solely on a sensitive attribute, a practice that may be unlawful in certain contexts. Furthermore, these methods can be advantageous in situations where applying distinct thresholds for different groups is impractical, or in cases where the decision-making body does not have access to the sensitive attribute. Such circumstances warrant the consideration of these methods despite their potential drawbacks in subgroup ranking. However, even in these cases, it is imperative to evaluate them at realistic positive rates.
An obvious limitation of our framework is that it operates under the assumption of having access to a static sensitive attribute, and it inherently faces difficulties in scenarios where these assumptions may not hold. For instance, the assumption of the sensitive attribute being static overlooks the evolving nature of certain attributes, such as gender, where people can identify as other categories over time. Another limitation arises when considering the assumption of unrestricted access to the sensitive attribute. In reality, legal and ethical considerations may impose constraints on obtaining or utilizing certain sensitive information (Haeri and Zweig, 2020; Veale and Binns, 2017; Johnson, 2021; Holstein et al., 2019). For instance, privacy laws and regulations may restrict the collection or use of specific attributes, further complicating the application of this framework in practice. These limitations underscore the need for ongoing refinement and adaptation of fairness frameworks to accommodate the complexities and nuances inherent in real-world scenarios.
Acknowledgments
This research was funded by Flemish Research Foundation (Grant number 11N7723N).
Appendix A Appendices
A.1 Score distributions for the other datasets
Figure 5: Score distributions for the Adult dataset
Figure 6: Score distributions for the Dutch dataset
Figure 7: Score distributions for the Law dataset
Figure 8: Score distributions for the Student dataset